Multiple Choice Question 81 A company expected its annual overhead costs to be $1500000 and direct labor costs to be $600000. Actual overhead was $1200000, and actual labor costs totaled $800000. How much is the companys predetermined overhead rate to the nearest cent? O $2.50 O $1.50 $1.88 O $2.00

Multiple Choice Question 97 Marigold Corp. applies overhead to production at a predetermined rate of 90% based on direct labor cost. Job No. 250, the only job still in process at the end of August, has been charged with manufacturing overhead of $6300. What was the amount of direct materials charged to Job 250 assuming the balance in Work in Process inventory is $15000? O $1700 6300. $7000. O $15000
Answer #1

Please provide rating...

Answer 1
Overhead rate = expected overhead cost / expected direct labor cost
=1500000/600000
         2.50
Therefore predetermined overhead rate = $2.50
Answer 2
Given that work in process inventory = 15000
Manufacturing overhead = 6300
Direct labor = 6300/90% 7000
Direct material = 15000-6300-7000 1700
Therefore correct answer is option $ 1700